OpenStudy (anonymous):

Volume of a sphere? I know the formula is V= 4/3 pi r^3 What is the volume of a sphere with the radius of 10^2 millimeters?

OpenStudy (bahrom7893):

plug in 10^2 instead of r and see what you get.

mathslover (mathslover):

Bahrom is right : \(V = \cfrac{4}{3} \pi r^3 \) Given to you that r = \(10^2\) mm So, \(V =\cfrac{4}{3} \pi (10^2)^3\)

mathslover (mathslover):

Don't forget the units too.

OpenStudy (anonymous):

Thanks guys!

mathslover (mathslover):

Candy, if you think Bahrom helped you, you can award him with a medal. Medal represents that he is a good helper and you acknowledge him! Click on the best response in front of his response. Thanks!